# Diffusion Models for Robotic Manipulation: A Survey

### v2

@@ 1 Introduction
Since 2022, there has been a noticeable increase in the implementation of diffusion probabilistic models within the field of robotic manipulation. These models are applied across various tasks, including trajectory planning, e.g., (Chi et al.,, 2023) and grasp prediction, e.g., (Urain et al.,, 2023). The ability of DMs to model multi-modal distributions is a great advantage in many robotic manipulation applications. In various manipulation tasks, such as trajectory planning and grasping, there exist multiple equally valid solutions (redundant solutions). Capturing all solutions improves generalizability and robots’ versatility, as it enables generating feasible solutions under different conditions, such as different placements of objects or different constraints during inference. Although in the context of trajectory planning using DMs, primarily imitation learning is applied, DMs have been adapted for integration with reinforcement learning (RL), e.g., (Geng et al.,, 2023). Research efforts focus on various components of the diffusion process adapted to different tasks in the domain of robotic manipulation. To give just some examples, developed architectures integrate different or even multiple input modalities. One example of an input modality could be point clouds (Ze et al.,, 2024; Ke et al.,, 2024). With the provided depth information, models can learn more complex tasks, for which a better 3D scene understanding is crucial. Another example of an additional input modality could be natural language (Ke et al.,, 2024; Du et al.,, 2023; Li et al.,, 2025), which also enables the integration of foundation models, like large language models, into the workflow. In Ze et al., (2024), both point clouds and language task instructions are used as multiple input modalities. Others integrate DMs into hierarchical planning (Ma et al., 2024b, ; Du et al.,, 2023) or skill learning (Liang et al.,, 2024; Mishra et al.,, 2023), to facilitate their state-of-the-art capabilities in modeling high-dimensional data and multi-modal distributions, for long-horizon and multi-task settings. Many methodologies, e.g. (Kasahara et al.,, 2024; Chen et al., 2023b, ), employ diffusion-based data augmentation in vision-based manipulation tasks to scale up datasets and reconstruct scenes. It is important to note that one of the major challenges of DMs is its comparatively slow sampling process, which has been addressed in many methods, e.g., (Song et al., 2021a, ; Chen et al.,, 2024; Zhou et al., 2024a, ), also enabling real-time prediction.
@@ 2.2.1 Improving Sampling Speed and Quality
There are also several methods (Zhou et al., 2024a, ; Li et al., 2024c, ; Wang et al., 2023c, ; Chen et al.,, 2024) regarding sampling speed, specifically for applications in robotic manipulation, which is different from the previously named methodologies, which were developed in the context of image processing. For example, Chen et al., (2024) samples from a more informed distribution than a Gaussian. They point out that even initial distributions approximated with simple heuristics result in better sample quality, especially when using few diffusion steps or when only a limited amount of data is available. Others (Prasad et al.,, 2024) use teacher–student distillation techniques (Tarvainen and Valpola,, 2017), where pretrained diffusion models serve as teachers, guiding student models to operate with larger denoising steps while preserving consistency with the teacher’s results at smaller steps. While this increases training effort, it decreases sampling time at inference, which is especially important in (near) real-time control.
@@ 4.1.1 Imitation Learning
In terms of the type of robotic embodiment, most works use parallel grippers or simpler end-effectors. However, few methods perform dexterous manipulation using DMs (Si et al.,, 2024; Ma et al., 2024a, ; Ze et al.,, 2024; Chen et al.,, 2024; Wang et al., 2024a, ; Freiberg et al.,, 2025; Welte and Rayyes,, 2025), to facilitate their stability and robustness, also in this high-dimensional setting.
@@
Classifier-free guidance Ho et al., (2021); Saha et al., (2024); Li et al., (2025); Power et al., (2023); Reuss et al., (2024, 2023)
@@
Reuss et al., (2024) RGB MV, Lan RHC ResNet, CLIP DiT
@@
Table 2: Technical details of trajectory diffusion using imitation learning. The references for the encoders are provided in Table 8.
In the following, the symbols and abbreviations are explained: H: Whether the method is hierarchical (✓) or not (✗). PCs: Point Clouds, Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, CT: complete trajectory in task space, J: complete trajectory in joint space. A “/” indicates that the information is not provided by the cited paper, while a “-” indicates that no specialized encoder is required as ground truth state information is used.
@@
As briefly discussed in Section 2.3, the entire trajectory can be generated as a single sample, multiple subsequences can be sampled using receding horizon control, or the trajectory can be generated by sampling individual steps. Only in a few methods (Janner et al.,, 2022; Ke et al.,, 2024) the whole trajectory is predicted at once. Although this enables a more efficient prediction, as the denoising has to be performed only once, it prohibits adapting to changes in the environment, requiring better foresight and making it unsuitable for more complex task settings with dynamic or open environments. On the other hand, sampling of individual steps increases the compounding error effect and can negatively affect temporal correlation. Instead of predicting micro-actions, some use DMs to predict waypoints (Shi et al.,, 2023). This can decrease the compounding error, by reducing the temporal horizon. However, it relies on preprocessing or task settings that ensure that the space in between waypoints is not occluded. Thus, typically, DMs generate trajectories consisting of sequences of micro-actions represented as end-effector positions, generally encompassing translation and rotation depending on end-effector actuation (Chi et al.,, 2023; Ze et al.,, 2024; Xu et al.,, 2023; Li et al., 2024c, ; Si et al.,, 2024; Scheikl et al.,, 2024; Ke et al.,, 2024; Ha et al.,, 2023). The control scheme is visualized in detail in Fig. 2. Although more commonly applied in grasp prediction, here the pose is sometimes also represented in special Euclidean group ($\mathbf{SE}(3)$) (Xian et al.,, 2023; Liu et al., 2023c, ; Ryu et al.,, 2024). Explained in more detail in Section 4.2, the group structure of the $\mathbf{SE}(3)$ Lie group enables continuous interpolation and transformations between multiple object poses. As (Liu et al., 2023c, ; Ryu et al.,, 2024) performs complex tasks involving trajectory planning and grasping for aligning multiple objects, these properties are important to ensure physically and geometrically grounded actions. However, as the prediction of $\mathbf{SE}(3)$ poses with DMs requires a more complex model structure and training in imitation learning, it is more usual to use representations, such as Euler angles or quaternions, in trajectory planning. Once the trajectory is sampled, the proximity of the predicted positions enables computing the motion between the positions with simple positional controllers without the need for complex trajectory planning techniques.
@@
As already discussed in Section 2.3 to ground the robots actions in the physical world, they are dependent on sensory input. Here, in the majority of methods, visual observations are used. While in the original work (Chi et al.,, 2023), combining visual robotic manipulation with DMs for trajectory planning, RGB-images are used, this does not provide sufficient geometrical information for intricate robotic tasks, especially in scenes containing occlusions. Thus, multiple later methods used 3D scene representations instead. Here, DMs are either directly conditioned on the point cloud (Li et al.,, 2025; Liu et al., 2023c, ; Wang et al., 2024a, ) or point cloud feature embeddings (Ze et al.,, 2024; Xian et al.,, 2023; Ke et al.,, 2024), from singleview (Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024a, ), or multiview camera setups (Ke et al.,, 2024; Xian et al.,, 2023). While multiview camera setups provide more complete scene information, they also require a more involved setup and more hardware resources.
@@
The ability of DMs to stably process high-dimensional input spaces enables the integration of multi-modal inputs, which is especially important in multi-skill tasks, to develop versatile and generalizable agents via arbitrary skill-chaining. Methodologies use videos (Xu et al.,, 2023), images, and natural language task instructions (Liang et al.,, 2024; Wang et al., 2024b, ; Zhou et al., 2024b, ), or even more diverse modalities, such as tactile information and point clouds (Wang et al., 2024b, ), to prompt skills.
@@
Another line of methods focuses on constrained trajectory learning. A typical goal is obstacle avoidance, object-centric, or goal-oriented trajectory planning, but other constraints can also be included. If the constraints are known prior to training, they can be integrated into the loss function. However, if the goal is to adhere to various and possibly changing constraint during inference another approach has to be taken. For less complex constraints, such as specific initial or goal states, (Janner et al.,, 2022) introduces a conditioning, where, after each denoising time step (Eq. 7), the particular state from the trajectory is replaced by the state from the constraint. However, this can lead the trajectory into regions of low likelihood, hence decreasing stability and potentially causing mode collapse. Moreover, this method is not applicable to more complex constraints.
One approach, also addressed by Janner et al., (2022), is classifier guidance (Dhariwal and Nichol,, 2021). Here, a separate model is trained to score the trajectory at each denoising step and steer it toward regions that satisfy the constraint. This is integrated into the denoising process by adding the gradient of the predicted score. It should be noted that for sequential data, such as trajectories, classifier guidance can also bias the sampling towards regions of low likelihood (Pearce et al.,, 2022). Thus, the weight of the guidance factor must be carefully chosen. Moreover, during the start of the denoising process the guidance model must predict the score on a highly uninformative output (close to Gaussian noise) and should have a lower impact. Therefore, it is important to inform the classifier of the denoising time step, train it also on noisy samples, or adjust the weight with which the guidance factor is integrated into the reverse process. Classifier guidance is applied in several methodologies (Mishra et al.,, 2023; Liang et al.,, 2023; Janner et al.,, 2022; Carvalho et al.,, 2023). However, it requires the additional training of a separate model. Furthermore, computing the gradient of the classifier at each sampling step adds additional computational cost. Thus, classifier-free guidance (Ho et al.,, 2021; Saha et al.,, 2024; Li et al.,, 2025; Power et al.,, 2023; Reuss et al.,, 2024, 2023) has been introduced, where a conditional and an unconditional DM per constraint are trained in parallel. During sampling, a weighted mixture of both DMs is used, allowing for arbitrary combinations of constraints, also not seen together during training. However, it does not generalize to entirely new constraints, as this would necessitate the training of new conditional DMs.
@@ 4.1.2 Offline Reinforcement Learning
Table 3: Technical details of trajectory diffusion using reinforcement learning. The references for the encoders are provided in Table 8.
In the following, the symbols and abbreviations are explained: H/S: Whether the method is hierarchical/skill-based (✓) or not (✗). Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), Eq.: Equivariant FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, Sia =single actions. A “-” indicates that no specialized encoder is required as ground truth state information is used.
@@ 4.2 Robotic grasp generation
Table 5: Technical details of grasp diffusion methodologies on $\mathbf{SE}(3)$ grasp synthesis. The references for the encoders are provided in Table 8. The references for the benchmarks are listed in Table 9.
In the following, the abbreviations used are explained: SDF: Signed Distance Function, TSDF: Truncated SDF, PCs: Point Clouds, FiLM: Convolutional Neural Network with Feature-wise Linear Modulation Perez et al., (2018), DiTs: Diffusion Transformers, Eq.: Equivariant, VN: Vector Neuron.
@@ 4.3.1 Scaling Data and Scene Augmentation
A challenge associated with data-driven approaches in robotics relates to substantial data requirements, which are time-consuming to acquire, particularly for real-world data. In the domain of imitation learning, it is essential to accumulate an adequate number of expert demonstrations that accurately represent the task at hand. While, by now, many methods, e.g. (Reuss et al.,, 2024; Ze et al.,, 2024; Ryu et al.,, 2024) only require a low number of five to fifty demonstrations, there are also methods, e.g. (Chen et al., 2023a, ; Saha et al.,, 2024) relying on more extensive data sets. Especially offline RL methods, e.g. (Carvalho et al.,, 2023; Ajay et al.,, 2023) usually require extensive amounts of data to accurately predict actions over the complete state-action space, also from suboptimal behavior. Moreover, increasing the variability in training data also has the potential to increase the generalizability of the learned policies. Thus, to automatically increase the variety and size of datasets, without additional costs on researchers and staff, or other more engineering-heavy autonomous data collection pipelines (Yu et al.,, 2023), many methodologies, e.g. (Chen et al., 2023b, ; Mandi et al.,, 2022), use DMs for data augmentation. In comparison to other strategies, such as domain randomization (Tremblay et al.,, 2018; Tobin et al.,, 2017), data augmentation with DMs directly augments the real-world data, making the data grounded in the physical world. In contrast, domain randomization requires complex tuning for each task, to ensure physical plausibility of the randomized scenes, and to enable sim-to-real transfer (Chen et al., 2023b, ).
@@
Given a set of real-world data, DM-based augmentation methods perform semantically meaningful augmentations via inpainting, such as changing object colors and textures (Zhang et al., 2024f, ), or even replacing whole objects, as well as corresponding language task descriptions (Chen et al., 2023b, ; Yu et al.,, 2023; Mandi et al.,, 2022). This enables both the augmentation of objects, which are part of the manipulation process, and backgrounds. The former increases the generalizability to different tasks and objects, while the latter increases robustness to scene information, which should not influence the policy. Some (Zhang et al., 2024f, ) also augment object positions and the corresponding trajectories to generate off-distribution demonstrations for DAgger, thus addressing the covariate shift problem in imitation learning. Others even (Katara et al.,, 2024) generate whole simulation scenes from given URDF files, prompted by a Large Language Model (LLM). Targeted towards offline RL methods, Di Palo et al., (2024) combines data augmentation with a form of hindsight-experience replay (Andrychowicz et al.,, 2017) to adapt the visual observations to the language-task instruction. This increases the number of successful executions in the replay buffer, which potentially increases the data efficiency. The method is used to learn policies for new tasks, on previously collected data, to align the data with the new task instructions.

### v3

@@ 1 Introduction
Since 2022, there has been a noticeable increase in the implementation of diffusion probabilistic models within the field of robotic manipulation. These models are applied across various tasks, including trajectory planning, e.g., (Chi et al.,, 2023) and grasp prediction, e.g., (Urain et al.,, 2023). The ability of DMs to model multi-modal distributions is a great advantage in many robotic manipulation applications. In various manipulation tasks, such as trajectory planning and grasping, there exist multiple equally valid solutions (redundant solutions). Capturing all solutions improves generalizability and robots’ versatility, as it enables generating feasible solutions under different conditions, such as different placements of objects or different constraints during inference. Although in the context of trajectory planning using DMs, primarily imitation learning is applied, DMs have been adapted for integration with reinforcement learning (RL), e.g., (Geng et al.,, 2023). Research efforts focus on various components of the diffusion process adapted to different tasks in the domain of robotic manipulation. To give just some examples, developed architectures integrate different or even multiple input modalities. One example of an input modality could be point clouds (Ze et al.,, 2024; Ke et al.,, 2024). With the provided depth information, models can learn more complex tasks, for which a better 3D scene understanding is crucial. Another example of an additional input modality could be natural language (Ke et al.,, 2024; Du et al.,, 2023; Li et al.,, 2025), which also enables the integration of foundation models, like large language models, into the workflow. In Ze et al., (2024), both point clouds and language task instructions are used as multiple input modalities. Others integrate DMs into hierarchical planning (Ma et al., 2024b, ; Du et al.,, 2023) or skill learning (Liang et al.,, 2024; Mishra et al.,, 2023), to facilitate their state-of-the-art capabilities in modeling high-dimensional data and multi-modal distributions, for long-horizon and multi-task settings. Many methodologies, e.g. (Kasahara et al.,, 2024; Chen et al., 2023b, ), employ diffusion-based data augmentation in vision-based manipulation tasks to scale up datasets and reconstruct scenes. It is important to note that one of the major challenges of DMs is its comparatively slow sampling process, which has been addressed in many methods, e.g., (Song et al., 2021a, ; Chen et al., 2024a, ; Zhou et al., 2024a, ), also enabling real-time prediction.
@@ 2.2.1 Improving Sampling Speed and Quality
There are also several methods (Zhou et al., 2024a, ; Li et al., 2024c, ; Wang et al., 2023c, ; Chen et al., 2024a, ) regarding sampling speed, specifically for applications in robotic manipulation, which is different from the previously named methodologies, which were developed in the context of image processing. For example, Chen et al., 2024a samples from a more informed distribution than a Gaussian. They point out that even initial distributions approximated with simple heuristics result in better sample quality, especially when using few diffusion steps or when only a limited amount of data is available. Others (Prasad et al.,, 2024) use teacher–student distillation techniques (Tarvainen and Valpola,, 2017), where pretrained diffusion models serve as teachers, guiding student models to operate with larger denoising steps while preserving consistency with the teacher’s results at smaller steps. While this increases training effort, it decreases sampling time at inference, which is especially important in (near) real-time control.
@@ 4.1.1 Imitation Learning
In terms of the type of robotic embodiment, most works use parallel grippers or simpler end-effectors. However, few methods perform dexterous manipulation using DMs (Si et al.,, 2024; Ma et al., 2024a, ; Ze et al.,, 2024; Chen et al., 2024a, ; Wang et al., 2024a, ; Freiberg et al.,, 2025; Welte and Rayyes,, 2025), to facilitate their stability and robustness, also in this high-dimensional setting.
@@
Classifier-free guidance Ho et al., (2021); Saha et al., (2024); Li et al., (2025); Power et al., (2023); Reuss et al., 2024a ; Reuss et al., (2023)
@@
Reuss et al., 2024a RGB MV, Lan RHC ResNet, CLIP DiT
@@
Reuss et al., 2024b RGB, Lan RHC ResNet, CLIP DiT
Table 2: Technical details of trajectory diffusion using imitation learning. The references for the encoders are provided in Table 8. In the following, the symbols and abbreviations are explained: H: Whether the method is hierarchical (✓) or not (✗). PCs: Point Clouds, Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, CT: complete trajectory in task space, J: complete trajectory in joint space. A “/” indicates that the information is not provided by the cited paper, while a “-” indicates that no specialized encoder is required as ground truth state information is used.
@@
As briefly discussed in Section 2.3, the entire trajectory can be generated as a single sample, multiple subsequences can be sampled using receding horizon control, or the trajectory can be generated by sampling individual steps. Only in a few methods (Janner et al.,, 2022; Ke et al.,, 2024) the whole trajectory is predicted at once. Although this enables a more efficient prediction, as the denoising has to be performed only once, it prohibits adapting to changes in the environment, requiring better foresight and making it unsuitable for more complex task settings with dynamic or open environments. On the other hand, sampling of individual steps increases the compounding error effect and can negatively affect temporal correlation. Instead of predicting micro-actions, some use DMs to predict waypoints (Shi et al.,, 2023). This can decrease the compounding error, by reducing the temporal horizon. However, it relies on preprocessing or task settings that ensure that the space in between waypoints is not occluded. Thus, typically, DMs generate trajectories consisting of sequences of micro-actions represented as end-effector positions, generally encompassing translation and rotation depending on end-effector actuation (Chi et al.,, 2023; Ze et al.,, 2024; Xu et al.,, 2023; Li et al., 2024c, ; Si et al.,, 2024; Scheikl et al.,, 2024; Ke et al.,, 2024; Ha et al.,, 2023). Once the trajectory is sampled, the proximity of the predicted positions enables computing the motion between the positions with simple positional controllers without the need for complex trajectory planning techniques. The control scheme is visualized in detail in Fig. 2. Although more commonly applied in grasp prediction, here the pose is sometimes also represented in special Euclidean group ($\mathbf{SE}(3)$) (Xian et al.,, 2023; Liu et al., 2023c, ; Ryu et al.,, 2024). Explained in more detail in Section 4.2, the group structure of the $\mathbf{SE}(3)$ Lie group enables continuous interpolation and transformations between multiple object poses. As (Liu et al., 2023c, ; Ryu et al.,, 2024) performs complex tasks involving trajectory planning and grasping for aligning multiple objects, these properties are important to ensure physically and geometrically grounded actions. However, as the prediction of $\mathbf{SE}(3)$ poses with DMs requires a more complex model structure and training in imitation learning, it is more usual to use representations, such as Euler angles or quaternions, in trajectory planning. Not only diffusion, but also flow matching has been adapted to use representations in  $\mathbf{SE}(3)$ or Riemannian manifolds in general (Braun et al.,, 2024).
@@
As already discussed in Section 2.3 to ground the robots actions in the physical world, they are dependent on sensory input. Here, in the majority of methods visual observations are used. In the original work (Chi et al.,, 2023), combining visual robotic manipulation with DMs for trajectory planning, the DM is conditioned on RGB-image observations. Many methods, e.g., (Si et al.,, 2024; Pearce et al.,, 2022; Li et al., 2024c, ), adopt using RGB inputs, also developing more intricate encoding schemes (Qi et al.,, 2025).
However, 2D visual scene representations may not provide sufficient geometrical information for intricate robotic tasks, especially in scenes containing occlusions. Thus, multiple later methods used 3D scene representations instead. Here, DMs are either directly conditioned on the point cloud (Li et al.,, 2025; Liu et al., 2023c, ; Wang et al., 2024a, ) or point cloud feature embeddings (Ze et al.,, 2024; Xian et al.,, 2023; Ke et al.,, 2024), from singleview (Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024a, ), or multiview camera setups (Ke et al.,, 2024; Xian et al.,, 2023). While multiview camera setups provide more complete scene information, they also require a more involved setup and more hardware resources.
@@
The ability of DMs to stably process high-dimensional input spaces enables the integration of multi-modal inputs, which is especially important in multi-skill tasks, to develop versatile and generalizable agents via arbitrary skill-chaining. Methodologies use videos (Xu et al.,, 2023), images, and natural language task instructions (Liang et al.,, 2024; Wang et al., 2024b, ; Zhou et al., 2024b, ; Reuss et al., 2024b, ), or even more diverse modalities, such as tactile information and point clouds (Wang et al., 2024b, ), to prompt skills.
@@
Another line of methods focuses on constrained trajectory learning. A typical goal is obstacle avoidance, object-centric, or goal-oriented trajectory planning, but other constraints can also be included. If the constraints are known prior to training, they can be integrated into the loss function. However, if the goal is to adhere to various and possibly changing constraints during inference, another approach has to be taken. For less complex constraints, such as specific initial or goal states, (Janner et al.,, 2022) introduces a conditioning, where, after each denoising time step (Eq. 7), the particular state from the trajectory is replaced by the state from the constraint. However, this can lead the trajectory into regions of low likelihood, hence decreasing stability and potentially causing mode collapse. Moreover, this method is not applicable to more complex constraints.
One approach, also addressed by Janner et al., (2022), is classifier guidance (Dhariwal and Nichol,, 2021). Here, a separate model is trained to score the trajectory at each denoising step and steer it toward regions that satisfy the constraint. This is integrated into the denoising process by adding the gradient of the predicted score. It should be noted that for sequential data, such as trajectories, classifier guidance can also bias the sampling towards regions of low likelihood (Pearce et al.,, 2022). Thus, the weight of the guidance factor must be carefully chosen. Moreover, during the start of the denoising process the guidance model must predict the score on a highly uninformative output (close to Gaussian noise) and should have a lower impact. Therefore, it is important to inform the classifier of the denoising time step, train it also on noisy samples, or adjust the weight with which the guidance factor is integrated into the reverse process. Classifier guidance is applied in several methodologies (Mishra et al.,, 2023; Liang et al.,, 2023; Janner et al.,, 2022; Carvalho et al.,, 2023). However, it requires the additional training of a separate model. Furthermore, computing the gradient of the classifier at each sampling step adds additional computational cost. Thus, classifier-free guidance (Ho et al.,, 2021; Saha et al.,, 2024; Li et al.,, 2025; Power et al.,, 2023; Reuss et al., 2024a, ; Reuss et al.,, 2023) has been introduced, where a conditional and an unconditional DM per constraint are trained in parallel. During sampling, a weighted mixture of both DMs is used, allowing for arbitrary combinations of constraints, also not seen together during training. However, it does not generalize to entirely new constraints, as this would necessitate the training of new conditional DMs.
@@ 4.1.2 Offline Reinforcement Learning
Table 3: Technical details of trajectory diffusion using reinforcement learning. The references for the encoders are provided in Table 8. In the following, the symbols and abbreviations are explained: H/S: Whether the method is hierarchical/skill-based (✓) or not (✗). Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), Eq.: Equivariant FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, Sia =single actions. A “-” indicates that no specialized encoder is required as ground truth state information is used.
@@ 4.2 Robotic grasp generation
Table 5: Technical details of grasp diffusion methodologies on $\mathbf{SE}(3)$ grasp synthesis. The references for the encoders are provided in Table 8. The references for the benchmarks are listed in Table 9. In the following, the abbreviations used are explained: SDF: Signed Distance Function, TSDF: Truncated SDF, PCs: Point Clouds, FiLM: Convolutional Neural Network with Feature-wise Linear Modulation Perez et al., (2018), DiTs: Diffusion Transformers, Eq.: Equivariant, VN: Vector Neuron.
@@ 4.3.1 Scaling Data and Scene Augmentation
A challenge associated with data-driven approaches in robotics relates to substantial data requirements, which are time-consuming to acquire, particularly for real-world data. In the domain of imitation learning, it is essential to accumulate an adequate number of expert demonstrations that accurately represent the task at hand. While, by now, many methods, e.g. (Reuss et al., 2024a, ; Ze et al.,, 2024; Ryu et al.,, 2024) only require a low number of five to fifty demonstrations, there are also methods, e.g. (Chen et al., 2023a, ; Saha et al.,, 2024) relying on more extensive data sets. Especially offline RL methods, e.g. (Carvalho et al.,, 2023; Ajay et al.,, 2023) usually require extensive amounts of data to accurately predict actions over the complete state-action space, also from suboptimal behavior. Moreover, increasing the variability in training data also has the potential to increase the generalizability of the learned policies. Thus, to automatically increase the variety and size of datasets, without additional costs on researchers and staff, or other more engineering-heavy autonomous data collection pipelines (Yu et al.,, 2023), many methodologies, e.g. (Chen et al., 2023b, ; Mandi et al.,, 2022), use DMs for data augmentation. In comparison to other strategies, such as domain randomization (Tremblay et al.,, 2018; Tobin et al.,, 2017), data augmentation with DMs directly augments the real-world data, making the data grounded in the physical world. In contrast, domain randomization requires complex tuning for each task, to ensure physical plausibility of the randomized scenes, and to enable sim-to-real transfer (Chen et al., 2023b, ).
@@
Given a set of real-world data, DM-based augmentation methods perform semantically meaningful augmentations via inpainting, such as changing object colors and textures (Zhang et al., 2024f, ), or even replacing whole objects, as well as corresponding language task descriptions (Chen et al., 2023b, ; Yu et al.,, 2023; Mandi et al.,, 2022). This enables both the augmentation of objects, which are part of the manipulation process, and backgrounds. The former increases the generalizability to different tasks and objects, while the latter increases robustness to scene information, which should not influence the policy. Some (Zhang et al., 2024f, ) also augment object positions and the corresponding trajectories to generate off-distribution demonstrations for DAgger, thus addressing the covariate shift problem in imitation learning. Others (Chen et al., 2024b, ) augment camera view, robot embodiments, or even (Katara et al.,, 2024) generate whole simulation scenes from given URDF files, prompted by a Large Language Model (LLM). Targeted towards offline RL methods, Di Palo et al., (2024) combines data augmentation with a form of hindsight-experience replay (Andrychowicz et al.,, 2017) to adapt the visual observations to the language-task instruction. This increases the number of successful executions in the replay buffer, which potentially increases the data efficiency. The method is used to learn policies for new tasks, on previously collected data, to align the data with the new task instructions.